A 500-kg satellite is in a circular orbit at an altitude of 500 km above the Earth’s surface. Because of air friction, the satellite eventually falls to the Earth’s surface, where it hits the ground with a speed of 2.00 km/s. How much energy was transformed into internal energy by means of air friction?




1
Expert's answer
2021-03-28T18:07:49-0400

"\text{Let}\newline
M=6*10^{24}kg\text{ Earth mass}"

"R = 6.4*10^7 m \text{ radius of the Earth}"

"G= 6.7*10^{-11}\text{ gravitational constant}"

"h = 0.5*10^7 m \text{ altitude of satellite }"

"m= 5*10^2\ kg\text{ satellite mass}"

"V\text{ satellite speed in orbit }"

"V_1=2*10^3m/s\text{ satellite landing speed}"

"\text{in orbit:}"

"F=ma"

"F= G\frac{mM}{(R+h)^2}"

"ma= G\frac{mM}{(R+h)^2}"

"a= G\frac{M}{(R+h)^2}"

"a=\frac{V^2}{(R+h)}"

"V^2= G\frac{M}{(R+h)}"

"E_k=\frac{mV^2}{2}=G\frac{mM}{2*(R+h)}"

"E_p=-G\frac{mM}{R+h}"

"E=E_k+E_p=-G\frac{mM}{2(R+h)}=-0.29*10^{10}"

"\text{upon landing:}"

"E_l=E_{p1}+E_{k1}"

"E_l=-G\frac{mM}{R}+\frac{mV_1^2}{2}=-0.30*10^{10}"

"Q= E-E_l=0.1*10^{10}"

Answer:"0.1*10^{10}J -\text{transformed into internal energy by means of air friction}"